# Continuous-time Risk-sensitive Reinforcement Learning via Quadratic Variation Penalty

### v1

@@ 5.1.1 Description of the algorithm
$-\frac{1}{T}\int_{0}^{T}\frac{\partial q}{\partial\psi_{2}^{-1}}\left(a_{i}(t);\boldsymbol{\psi}_{i}\right)\left[\mathrm{d}V\left(t,X_{i}(t);\theta_{i}\right)-q\left(a_{i}(t);\boldsymbol{\psi}_{i}\right)\mathrm{d}t+\frac{1-\gamma}{2}\mathrm{d}\langle V\left(\cdot,X_{i}(\cdot);\theta_{i}\right)\rangle\right],$ (25)
@@
The full iterative procedure is summarized by Algorithm 3.
@@ 5.1.2 Convergence of the algorithm
The analysis of this algorithm falls into the scope of the general stochastic approximation algorithms [^8] The development of stochastic approximation can be traced back to Robbins and Monro, (1951) and Lai, (2003). More introduction can be found, e.g., in Kushner and Yin, (2003).. The implementation in Algorithm 3 essentially suggests that the time discretization is only necessary for computing integrals in (23), (24), and (25) whose error is governed by the standard numerical analysis. Hence our interest is mainly to analyze the expectation and variance (conditioned on $\theta_{i},\boldsymbol{\psi}_{i}$) of (23), (24), and (25), respectively, as in stochastic approximation algorithms. However, in our examples, the variance of these increment is not uniformly bounded, but growing. Hence the projection is necessary to ensure the variance does not grow too fast. The ideas of projection are borrowed from Andradóttir, (1995), and the convergence rate for the recursion is borrowed from Broadie et al., (2011).
@@
Suppose that the temperature parameter $\lambda$ is a positive constant along the learning procedure (26), and positive sequences $\{a_{\psi,i}\}_{i\geq 0}$, $\{b_{i}\}_{i\geq 0},\{c_{i}\}_{i\geq 0}$ satisfy
@@
    $b_{i}\uparrow\infty,c_{i}\uparrow\infty$;
@@
    $\sum_{i=1}^{\infty}a_{\psi,i}^{2}b_{i}^{2}c_{i}^{4}<\infty$.
Then the learning procedure described by (26) will converge almost surely. Moreover, with suitable choices of $a_{\psi,n}\sim\frac{1}{n}$, $b_{n},c_{n}\sim\log n$, the mean squared error (MSE) of the parameters in the policy converges to zero at the rate $\mathbb{E}[|\boldsymbol{\psi}_{n}-\boldsymbol{\psi}^{*}|^{2}]=\tilde{O}(\frac{1}{n})$, where $\tilde{O}(\cdot)$ means the big O ignoring the logarithm factor.
@@
The next theorem characterizes how fast the accumulated ERWL up to the first $N$ episodes grows in $N$, depending on whether a deterministic policy can be executed. If ERWL is diminishing, then the accumulated ERWL will grow in $N$ at a sublinear rate. The slower it grows, the more efficient the algorithm is.
@@
    If the deterministic policy $\mathcal{N}(\psi_{1,i},0)$ is executed along the learning procedure (26), then with suitable choices of sequences $a_{\psi,n}\sim\frac{1}{n}$, $b_{n},c_{n}\sim\log n$, $\lambda_{n}=\lambda$, the expected accumulated ERWL satisfies
@@
    If the stochastic policy $\mathcal{N}(\psi_{i,1},\lambda_{i}\psi_{i,2})$ is executed along the learning procedure (26), then with suitable choices of sequences $a_{\psi,n}\sim\frac{1}{\sqrt{n}}$, $b_{n}=O(1)$, $c_{n}\sim\log n$, $\lambda_{n}\sim\frac{1}{\sqrt{n}}$, the expected accumulated ERWL satisfies
@@ 5.1.3 Numerical results
I use two sets of tuning parameters that correspond to two situations in Theorem 5 to illustrate. In the first setting, the temperature parameter is fixed as $\lambda=3$, and the step size $a_{\psi,n}$ decays as $n^{-1}$ and $b_{n},c_{n}$ grows at the rate $\log n$. In the second setting, the temperature parameter decays as $\lambda n^{-1/2}$, and the step size $a_{\psi,n}$ decays as $n^{-1/2}$, $b_{n}$ is a fixed small constant, and $c_{n}$ grows at the rate $\log n$. I repeat the simulation runs for 1000 times to estimate the mean squared error or the mean equivalent relative wealth of the learning algorithm.
@@
Figure 1 illustrates the performance of the algorithm under the first tuning parameters setup. This setup satisfies the conditions in Theorem 4 and Theorem 5 (a). The left panel shows the convergence rate of MSE almost match the theoretical results. The accumulated ERWL on the right panel (in the log-log scale) does not grow linearly.
@@
Figure 2 illustrates the performance of the algorithm under the first tuning parameters setup. This setup satisfies the conditions in Theorem 5 (b). It is interesting to notice that here that the MSE of $\psi_{1,n}-\psi_{1}^{*}$ also decays almost at the rate of $n^{-1}$, however, in the proof, this rate is only $\tilde{O}(n^{-1/2})$. [^9] Proving a faster convergence rate of MSE is possible by a more careful analysis on the variance of the increment term, which would also decays in $\lambda_{n}^{2}$. A faster convergence rate in the mean of the learned policy cannot further improve the order of the accumulated ERWL because the variance in the learned policy cannot be further reduced.
(a) The mean squared error of the learned parameters in the policy.
(b) The mean accumulated ERWL of the learned deterministic policy.
@@
Figure 1: The illustration of the learned policy. The left panel shows the convergence of the mean squared error of the learned parameters in the policy, and the right panel shows the mean equivalent relative wealth loss of the learned deterministic policy. Both panels are in the log-scales. The results are based on simulated data with 1000 runs. The shaded area indicates twice the standard deviation of the estimated expectation. The temperature parameter is taken as $\lambda=3$, and the learning rate $a_{\psi,n}=(1+n)^{-1}$. The number of episode within each simulation run is $10^{5}$.
(a) The mean squared error of the learned parameters in the policy.
(b) The mean accumulated ERWL of the learned stochastic policy.
@@
Figure 2: The illustration of the learned policy. The left panel shows the convergence of the mean squared error of the learned parameters in the policy, and the right panel shows the mean equivalent relative wealth loss of the learned deterministic policy. Both panels are in the log-scales. The results are based on simulated data with 1000 runs. The shaded area indicates twice the standard deviation of the estimated expectation. The temperature parameter is taken as $\lambda_{n}=3(n+1)^{-1/2}$, and the learning rate $a_{\psi,n}=(n+1)^{-1/2}$. The number of episode within each simulation run is $10^{5}$.
@@ 5.2.1 Description of the algorithm
$V(X;\boldsymbol{\theta})=$ $\theta_{2}x^{2}+\theta_{1}x,$
@@ 5.2.2 Numerical results
In the numerical experiments, the configurations are $A=-2$, $B=1$, $C=0.25$, $D=1$, $M=N=Q=2$, $R=P=1$. The behavior policy is taken as state-independent, i.e., $\boldsymbol{\pi}^{b}(\cdot|X)=\mathcal{N}(0,1)$, and the initial state is $x_{0}=0$. The time discretization for the sample is $\Delta t=0.01$. The sample reward is generated as $\mathcal{R}|a,X\sim\mathcal{N}\left(r(X,a),\Delta t\right)$. I mainly examine the effect of the sample size and the risk sensitivity coefficient. Hence, I fix the temperature parameter $\lambda=2$, and generate sample with length $T\in\{1,10,100\}$ and choose the risk sensitivity coefficient $\epsilon\in\{0,-0.1,-0.5,-1,-2,-5,-10,-20\}$. In each simulation run, I use the behavior policy to generate sample with length $T$, and then conduct learning algorithms to obtain the learned parameters $\hat{\boldsymbol{\psi}},\hat{\boldsymbol{\theta}}$, and finally compute its distance to the optimal policy to the problem (29). The simulation runs are repeated for $10^{4}$ times to evaluate the mean squared error (MSE).
@@
The results are presented in Table 1. The top, middle, bottom panel stands for different sample sizes. In each panel, notice that solving the non-risk-sensitive objective does not lead to the most accurate estimate for the policy. With suitable values of $\epsilon$, solving the risk-sensitive objective reduces MSE. As $\epsilon$ approaches 0, the results get closer to the non-risk-sensitive objective, whereas overly large $\epsilon$ (in the absolute value) harms the accuracy. Overall, as the sample size increases, the performance get significantly better. The optimal $\epsilon$ is around $-5$ when the sample size is small, and is around $-2$ when the sample size is large. It confirms the intuition that large sample size provides a better approximation to the ground truth distribution, and hence the agent does not have to be too risk sensitive.
@@
$\epsilon$ 00 $-0.1$ $-0.5$ $-1$ $-2$ $-5$ $-10$ $-20$
$T=1$ MSE of $\hat{\psi}_{1}$ 1.5874 1.5560 1.4129 1.2482 1.0099 0.7923 0.8638 0.9614
MSE of $\hat{\psi}_{2}$ 0.3651 0.3984 0.2736 0.2413 0.2417 0.3384 0.5560 0.7940
MSE of $\hat{\psi}_{3}$ 0.3980 0.4183 0.1968 0.1215 0.1141 0.1566 0.2705 0.3845
$T=10$ MSE of $\hat{\psi}_{1}$ 0.1119 0.1003 0.0691 0.0640 0.1129 0.6625 1.8854 3.8925
MSE of $\hat{\psi}_{2}$ 0.0341 0.0318 0.0266 0.0334 0.0957 0.6144 1.8258 3.8690
MSE of $\hat{\psi}_{3}$ 0.0130 0.0064 0.0056 0.0055 0.0492 0.3054 0.8281 1.7897
$T=100$ MSE of $\hat{\psi}_{1}$ 0.0083 0.0081 0.0076 0.0071 0.0064 0.0714 0.9555 5.7822
MSE of $\hat{\psi}_{2}$ 0.0043 0.0042 0.0039 0.0037 0.0041 0.0675 0.9008 5.7014
MSE of $\hat{\psi}_{3}$ 0.0003 0.0003 0.0003 0.0004 0.0006 0.0209 0.3404 2.2558
6 Conclusion
In this paper, I investigated the continuous-time risk-sensitive RL problems in the exponential form. This formulation is linked to the mean-variance problems when the risk sensitivity coefficient is close to zero, and is a relaxed formulation of the distributional robust RL when the uncertainty set is specified using KL divergence. Using the properties of the exponential martingale, I turned the risk-sensitive objective into a usual additive objective plus an extra penalty term on the QV of the value-to-go process. Thus, it becomes a problem to solve a recursive utility maximization problem. Such an equivalent formulation exists in the classical risk-sensitive control problems and also applies to the entropy-regularized RL problems. The benefit of such transformation avoids the multiplicative Bellman equation analyzed in previous risk-sensitive MDP literature, and numerically avoids the computation of exponential of a possibly large number.
I defined the q-function in this context which differs from the q-function in the non-risk-sensitive problems (Jia and Zhou, , 2023) by an extra term. However, the optimal q-function and the optimal value function can still be jointly pinned down using the martingale characterization, which also involves in the QV of the value function process. It also extends to ergodic tasks under stronger regularity conditions. Interestingly, I point out that the connection between the q-learning and the policy gradient in the non-risk-sensitive problems does not hold anymore in risk-sensitive problems, because the QV penalty is inherently a nonlinear term in the value function.
I demonstrate the proposed algorithm on two applications. For Merton’s investment problem with power utility functions, it can be embedded to a risk-sensitive objective in terms of the log returns. When the stock price follows the Black-Scholes model, I provide theoretical analysis on the proposed algorithm to show its convergence and the convergence rate. In the off-policy learning for a linear-quadratic problem, numerically, I illustrate that when the sample size is limited, incorporating the risk-sensitive objective to conduct learning indeed improves the learning accuracy compared to that under the non-risk-sensitive objective. The best choice of the risk sensitivity coefficient falls in the sample size. This result is consistent with the relation between the risk-sensitive RL and the distributional robust RL, in the sense that including extra risk sensitivity concerns can address the discrepancy between a finite data set and its population distribution.
The analysis of RL algorithm in this paper is restricted to simple cases, but it sheds light on the role of the temperature and the entropy regularization play in the learning procedure from the algorithmic perspective. In Merton’s investment problem analyzed in this paper, I show that the product between the temperature parameter and the step size in the iteration is equivalent to an “effective learning rate” under the stochastic approximation framework. On one hand, low temperature slows down the equivalent learning rate, and in particular, zero temperature implies that there are no effective learning signals to update the q-function. On the other hand, low temperature induces less randomness in the policy, and hence, overall reduces the noises in the learning signals and leverages more on the learned optimal decisions. Therefore, the tuning of temperature parameter reflects the tradeoff between exploration (speed of learning) and exploitation. When the temperature parameter and the step size decay at compatible rates, the desired convergence, convergence rate, and regret (in terms of equivalent relative wealth loss) can be guaranteed.
The risk-sensitive objective in the exponential form considered in this paper is one special case of general functional classes, and the associated KL divergence is one tractable metric to reflect the ambiguity in the distributions. How to address the distributional robustness other than the KL divergence in the continuous-time diffusion processes in a tractable way and to conduct RL still remains largely unclear. In addition, alternative formulation to address risk sensitivity may lack time-consistency, and dynamic programming-based RL methods cannot be applied therein. Even within the continuous-time risk-sensitive RL framework presented in this paper, there are still many theoretical and practical questions open for future research. For example, the analysis and applications of q-learning algorithms in general cases, and the comparison of generalization error between risk-sensitive objective and non-risk-sensitive objective in the offline, off-policy settings.

### v2

@@ 5.1.1 Description of the algorithm
$\frac{1}{T}\int_{0}^{T}\frac{\partial q}{\partial\psi_{2}^{-1}}\left(a_{i}(t);\boldsymbol{\psi}_{i}\right)\left[\mathrm{d}V\left(t,X_{i}(t);\theta_{i}\right)-q\left(a_{i}(t);\boldsymbol{\psi}_{i}\right)\mathrm{d}t+\frac{1-\gamma}{2}\mathrm{d}\langle V\left(\cdot,X_{i}(\cdot);\theta_{i}\right)\rangle\right],$ (25)
@@
The full iterative procedure is summarized by Algorithm 3 by discretely sampling the action sequences and using finite sums to approximate the integrals in (23), (24), and (25).
@@ 5.1.2 Convergence of the algorithm
The analysis of this algorithm falls into the scope of the general stochastic approximation algorithms (cf. Kushner and Yin, (2003)). The implementation in Algorithm 3 essentially suggests that the time discretization is only necessary for computing integrals in (23), (24), and (25) whose error is governed by the standard numerical analysis. Hence our interest is mainly to analyze the expectation and variance (conditioned on $\theta_{i},\boldsymbol{\psi}_{i}$) of (23), (24), and (25), respectively, as in stochastic approximation algorithms. However, in our examples, the variance of these increments is not uniformly bounded, but growing. Hence the projection is necessary to ensure the variance does not grow too fast. The ideas of projection are borrowed from Andradóttir, (1995), and the convergence rate for the recursion is borrowed from Broadie et al., (2011).
@@
Suppose that the temperature parameter $\lambda$ is a positive constant along the learning procedure (26), with time discretization step size $\{\Delta t_{i}\}_{i\geq 0}$. Suppose there are positive sequences $\{a_{\psi,i}\}_{i\geq 0}$, $\{b_{i}\}_{i\geq 0},\{c_{i}\}_{i\geq 0}$ satisfying
@@
    $b_{i}\uparrow\infty,c_{i}\uparrow\infty,\Delta t_{i}\downarrow 0$;
@@
    $\sum_{i=1}^{\infty}(a_{\psi,i}^{2}b_{i}^{4}c_{i}^{8}+a_{\psi,i}\Delta t_{i}c_{i}^{3})<\infty$.
[…]
@@
The next theorem characterizes how fast the accumulated ERWL up to the first $N$ episodes grows in $N$, depending on whether a deterministic policy can be executed. If ERWL is diminishing, then the accumulated ERWL will grow in $N$ at a sublinear rate. The slower it grows, the more efficient the algorithm is. In particular, I demonstrate the regret that concerns the learned deterministic policy $\mathcal{N}(\psi_{1,i},0)$ as well as the learned stochastic policy $\mathcal{N}(\psi_{i,1},\lambda_{i}\psi_{i,2})$. Note that compared to the deterministic policy, a stochastic policy incurs more losses because it lacks exploitation, which results in larger regret. Here, two metrics concern the same learning algorithm, which has to rely on the stochastic policy in the learning process to ensure the convergence. Moreover, Theorem 5 also illustrates the different requirements on tuning the learning rate and temperature sequences to achieve the desired regret orders.
@@
    If the deterministic policy $\mathcal{N}(\psi_{1,i},0)$ is executed along the learning procedure (26), then with suitable choices of sequences $a_{\psi,n},\Delta t_{n}\sim\frac{1}{n}$, $b_{n},c_{n}\sim\log n$, $\lambda_{n}=\lambda$, the expected accumulated ERWL satisfies
@@
    If the stochastic policy $\mathcal{N}(\psi_{i,1},\lambda_{i}\psi_{i,2})$ is executed along the learning procedure (26), then with suitable choices of sequences $a_{\psi,n},\Delta t_{n}\sim\frac{1}{\sqrt{n}}$, $b_{n}=O(1)$, $c_{n}\sim\log n$, $\lambda_{n}\sim\frac{1}{\sqrt{n}}$, the expected accumulated ERWL satisfies
@@ 5.1.3 Numerical results
I use two sets of tuning parameters that correspond to two situations in Theorem 5 to illustrate. In the first setting, the temperature parameter is a fixed constant, and we take it as $\lambda=3,1,0.3$, and the step size $a_{\psi,n}$ decays as $n^{-1}$ and $b_{n},c_{n}$ grows at the rate $\log n$. In the second setting, the temperature parameter decays as $\lambda n^{-1/2}$, and the step size $a_{\psi,n}$ decays as $n^{-1/2}$, $b_{n}$ is a fixed small constant, and $c_{n}$ grows at the rate $\log n$. I repeat the simulation runs for 1000 times to estimate the mean squared error or the mean equivalent relative wealth of the learning algorithm.
@@
Figure 1 illustrates the performance of the algorithm under the first tuning parameters setup. This setup satisfies the conditions in Theorem 4 and Theorem 5 (a). The left two panels show the convergence rate of MSE almost match the theoretical results. The accumulated ERWL on the right panel (in the log-log scale) does not grow linearly. In particular, when $\lambda$ is too small ($\lambda=0.3$), the algorithm incurs larger error in the long run; whereas, the rate of convergence of the mean squared error of the learned $\psi_{1,n}$ is the same. The reason that the mean squared error of the learned $\psi_{2,n}$ does not exhibit the same convergence rate when $\lambda=0.3$ is either because the initial learning rate is not large enough or the number of episodes is still large enough to visualize the order.
@@
Figure 2 illustrates the performance of the algorithm under the first tuning parameters setup. This setup satisfies the conditions in Theorem 5 (b). It is interesting to notice that here that the MSE of $\psi_{1,n}-\psi_{1}^{*}$ also decays almost at the rate of $n^{-1}$, however, in the proof, this rate is only $\tilde{O}(n^{-1/2})$. [^9] Proving a faster convergence rate of MSE is possible by a more careful analysis on the variance of the increment term, which would also decays in $\lambda_{n}^{2}$. A faster convergence rate in the mean of the learned policy cannot further improve the order of the accumulated ERWL because the variance in the learned policy cannot be further reduced. When the temperature parameter also decays in the number of episode, the value of initial temperature also has similar impact as in Figure 1. Moreover, in this case, the accumulated ERWL when initial temperature is lower also tends to be smaller. This is because in this case, ERWL is calculated for the stochastic policy; and lower temperature implies lower variance in the policy, and hence improves its performance.
(a) The mean squared error of the learned $\psi_{1}$.
(b) The mean squared error of the learned $\psi_{2}$.
(c) The mean accumulated ERWL of $\mathcal{N}(\psi_{1,n},0)$.
@@
Figure 1: The illustration of the learned policy. The left two panels show the convergence of the mean squared error of the learned parameters in the policy, and the right panel shows the mean equivalent relative wealth loss of the learned deterministic policy. Both panels are in the log-scales. The results are based on simulated data with 1000 runs. The shaded area indicates twice the standard deviation of the estimated expectation. The temperature parameters are taken as $\lambda=3,1,0.3$, respectively, and the learning rate $a_{\psi,n}=10/(1+n)$. The number of episodes within each simulation run is $10^{5}$.
(a) The mean squared error of the learned $\psi_{1}$.
(b) The mean squared error of the learned $\psi_{2}$.
(c) The mean accumulated ERWL of $\mathcal{N}(\psi_{1,n},\lambda_{n}\psi_{2,n})$.
@@
Figure 2: The illustration of the learned policy. The left two panels show the convergence of the mean squared error of the learned parameters in the policy, and the right panel shows the mean equivalent relative wealth loss of the learned deterministic policy. Both panels are in the log-scales. The results are based on simulated data with 1000 runs. The shaded area indicates twice the standard deviation of the estimated expectation. The temperature parameter is taken as $\lambda_{n}=3/\sqrt{n+1},1/\sqrt{n+1},0.3/\sqrt{n+1}$, respectively, and the learning rate $a_{\psi,n}=10/\sqrt{n+1}$. The number of episodes within each simulation run is $10^{5}$.
@@ 5.2.1 Description of the algorithm
$V(X;\boldsymbol{\theta})=$ $\theta_{2}x^{2}+\theta_{1}x.$
@@ 5.2.2 Numerical results
In the numerical experiments, the configurations are A=2A=-2, $B=1$, $C=0.25$, $D=1$, $M=N=Q=2$, $R=P=1$. The behavior policy is taken as state-independent, i.e., $\boldsymbol{\pi}^{b}(\cdot|X)=\mathcal{N}(0,1)$, and the initial state is $x_{0}=0$. The time discretization for the sample is $\Delta t=0.01$. The sample reward is generated as $\mathcal{R}|a,X\sim\mathcal{N}\left(r(X,a),\Delta t\right)$. I mainly examine the effect of the sample size and the risk sensitivity coefficient. Hence, I choose the temperature parameter $\lambda\in\{3,1,0.3\}$, and generate sample with length $T\in\{1,10,100\}$ and choose the risk sensitivity coefficient $\epsilon\in\{0,-0.1,-0.5,-1,-2,-5,-10,-20\}$. In each simulation run, I use the behavior policy to generate sample with length $T$, and then conduct learning algorithms to obtain the learned parameters $\hat{\boldsymbol{\psi}},\hat{\boldsymbol{\theta}}$, and finally compute its distance to the optimal policy to the problem (29). The simulation runs are repeated for $10^{4}$ times to evaluate the mean squared error (MSE).
@@
The results are presented in Table 1. The top, middle, bottom panel stands for different sample sizes, and subtables (a), (b) and (c) correspond to different temperature parameters. In each panel, notice that solving the non-risk-sensitive objective does not lead to the most accurate estimate for the policy. With suitable values of $\epsilon$, solving the risk-sensitive objective reduces MSE. As $\epsilon$ approaches 0, the results get closer to the non-risk-sensitive objective, whereas overly large $\epsilon$ (in the absolute value) harms the accuracy. Overall, as the sample size increases, the performance get significantly better. The optimal $\epsilon$ is between 5-5 and 2-2 when the sample size is small, and is between 2-2 and 1-1 when the sample size is large. It confirms the intuition that large sample size provides a better approximation to the ground truth distribution, and hence the agent does not have to be too risk sensitive. In addition, a smaller temperature parameter $\lambda$ seems to induce an overall smaller estimation error in the off-policy, offline setting. This is probably because, unlike the on-policy, online learning, in this setting, a data set has been generated and fixed, and thus, the benefits of exploration are limited; instead, a problem with larger temperature deviates more from the original objective function.
$\epsilon$ 0 0.1-0.1 0.5-0.5 1-1 2-2 5-5 10-10 20-20
$T=1$ MSE of $\hat{\psi}_{1}$ 2.3731 2.3218 2.1051 1.7744 1.3750 0.9908 0.8914 1.0191
MSE of $\hat{\psi}_{2}$ 0.3960 0.4287 0.3055 0.2514 0.2383 0.3082 0.4971 0.7456
MSE of $\hat{\psi}_{3}$ 0.4760 0.4631 0.2284 0.1211 0.1007 0.1243 0.2338 0.3388
$T=10$ MSE of $\hat{\psi}_{1}$ 0.3077 0.2696 0.1473 0.1057 0.1272 0.4491 1.2589 2.9190
MSE of $\hat{\psi}_{2}$ 0.0550 0.0529 0.0294 0.0322 0.0790 0.3864 1.1444 2.6676
MSE of $\hat{\psi}_{3}$ 0.0385 0.0200 0.0043 0.0098 0.0305 0.1922 0.6222 1.4192
$T=100$ MSE of $\hat{\psi}_{1}$ 0.0186 0.0178 0.0156 0.0138 0.0116 0.0360 0.4659 3.2283
MSE of $\hat{\psi}_{2}$ 0.0067 0.0061 0.0047 0.0037 0.0031 0.0322 0.4406 2.9502
MSE of $\hat{\psi}_{3}$ 0.0004 0.0004 0.0003 0.0004 0.0005 0.0151 0.2280 1.3929
(a) When the temperature parameter $\lambda=3$.
$\epsilon$ 0 0.1-0.1 0.5-0.5 1-1 2-2 5-5 10-10 20-20
$T=1$ MSE of $\hat{\psi}_{1}$ 0.923 0.913 0.905 0.832 0.801 1.003 1.486 2.336
MSE of $\hat{\psi}_{2}$ 0.332 0.350 0.292 0.324 0.426 0.777 1.356 2.266
MSE of $\hat{\psi}_{3}$ 0.313 0.326 0.201 0.153 0.170 0.328 0.607 1.113
$T=10$ MSE of $\hat{\psi}_{1}$ 0.038 0.037 0.036 0.038 0.075 0.287 1.065 2.876
MSE of $\hat{\psi}_{2}$ 0.012 0.012 0.012 0.014 0.044 0.258 1.063 3.053
MSE of $\hat{\psi}_{3}$ 0.003 0.003 0.002 0.003 0.013 0.090 0.328 1.121
$T=100$ MSE of $\hat{\psi}_{1}$ 0.0042 0.0041 0.0040 0.0038 0.0039 0.2119 2.2732 10.8350
MSE of $\hat{\psi}_{2}$ 0.0027 0.0027 0.0027 0.0029 0.0045 0.2200 2.4270 12.3801
MSE of $\hat{\psi}_{3}$ 0.0002 0.0003 0.0003 0.0004 0.0008 0.0565 0.7789 4.5567
(b) When the temperature parameter $\lambda=1$.
$\epsilon$ 0 0.1-0.1 0.5-0.5 1-1 2-2 5-5 10-10 20-20
$T=1$ MSE of $\hat{\psi}_{1}$ 0.2989 0.2989 0.2988 0.2988 0.2987 0.3097 0.3788 0.5657
MSE of $\hat{\psi}_{2}$ 0.0491 0.0492 0.0494 0.0498 0.0505 0.0668 0.1329 0.3346
MSE of $\hat{\psi}_{3}$ 0.0556 0.0556 0.0556 0.0556 0.0556 0.0604 0.0853 0.1520
$T=10$ MSE of $\hat{\psi}_{1}$ 0.2071 0.2070 0.2066 0.2061 0.2051 0.2036 0.2299 0.3840
MSE of $\hat{\psi}_{2}$ 0.0172 0.0172 0.0172 0.0173 0.0175 0.0191 0.0439 0.2008
MSE of $\hat{\psi}_{3}$ 0.0299 0.0298 0.0297 0.0295 0.0292 0.0287 0.0289 0.0612
$T=100$ MSE of $\hat{\psi}_{1}$ 0.0031 0.0031 0.0030 0.0029 0.0050 0.3170 4.1180 15.1154
MSE of $\hat{\psi}_{2}$ 0.0023 0.0023 0.0023 0.0026 0.0066 0.4215 4.7278 17.6495
MSE of $\hat{\psi}_{3}$ 0.0002 0.0002 0.0003 0.0004 0.0015 0.0963 1.1864 6.2771
(c) When the temperature parameter $\lambda=0.3$.
@@
6 Discussions
In this paper, I studied continuous-time risk-sensitive RL in the exponential form. By leveraging exponential martingale properties, I transformed the risk-sensitive objective into a standard additive form with an additional quadratic variation penalty. I introduced the q-function in this setting and showed that the optimal q-function and value function can still be jointly characterized through a linear martingale approach. However, unlike in the risk-neutral case, the usual connection between q-learning and policy gradient breaks down due to the inherent nonlinearity of the value function.
In contrast to discrete-time risk-sensitive problems, where an equivalent simplification is lacking and algorithms tend to be more complex, the continuous-time formulation offers significant advantages. The key difference lies in the treatment of transitions: in discrete time, the transition probabilities remain unknown and difficult to learn, leading to nonlinear recursions and the challenging multiplicative Bellman equation. In contrast, the continuous-time setting allows for a transformation that avoids these complications both theoretically and numerically. Theoretically, it replaces the multiplicative Bellman equation with a simpler additive structure, and numerically, it eliminates the need to compute exponentials of potentially large values. This advantage stems from the fact that in continuous time, a normal approximation holds incrementally, enabling a more tractable and efficient learning process.
I demonstrate the proposed algorithm on two applications: Merton’s investment problem with power utility functions under the Black-Scholes model, and the off-policy learning for a linear-quadratic problem. These two serve as simple test cases where the desired solutions are known to take linear or even constant forms. However, real-world problems often involve complex, unknown q-functions and policies, requiring function approximation methods such as neural networks. This introduces approximation errors that can impact performance. Moreover, the complicated functional forms may lead to non-convex optimization landscape, increasing numerical instability. These challenges, which are not captured in my experiments, highlight the need for further research to assess the algorithm’s robustness and scalability in more complex, high-dimensional settings.
The analysis of RL algorithm in this paper is restricted to simple cases, but it provides key insights into the role of the temperature $\lambda$ and the entropy regularization in the learning procedure. In the context of Merton’s investment problem, I show that the product between the temperature parameter and the step size acts as an “effective learning rate” under the stochastic approximation framework. This reveals a fundamental tradeoff: while lower temperatures slow down learning and, in the extreme case of zero temperature, eliminate effective learning signals, they also reduce randomness in the policy, leading to less noisy updates. Moreover, when the temperature and step size decay at compatible rates, the algorithm achieves the desired convergence rate. This analysis serves as a first step in understanding the impact of the temperature parameter on learning algorithms while explicitly accounting for sampling error – an aspect that appears to be overlooked in prior literature. The approach in this paper sets an example for proving the convergence of stochastic-approximation-based algorithms for continuous-time, continuous-space RL. However, extending this methodology to more complex settings remains challenging, because verifying the typical convergence conditions in general cases is nontrivial. These open questions highlight important directions for future research.
The numerical study on off-policy linear-quadratic problems provides valuable insights into the role of the risk sensitivity coefficient $\epsilon$. The results suggest that when the sample size is limited, incorporating a risk-sensitive objective can improve learning accuracy compared to a risk-neutral approach. Moreover, the optimal choice of $\epsilon$ depends on the sample size, as reflected in the connection between risk-sensitive RL and distributionally robust RL. Specifically, introducing risk sensitivity helps mitigate the discrepancy between a finite data set and the underlying population distribution. As a direction for future work, it would be valuable to establish theoretical guarantees on how risk-sensitive formulations address distributional shift and to develop data-driven methods for selecting the risk sensitivity coefficient.